# Simple Sufficient Criteria for Optimality of Entanglement Witnesses

### v1

@@ I Introduction
In this work we tackle both these problems. As such, this paper is organized as follows: In Sec. II.1 we recap some basic notions of channels and general linear maps, and Sec. II.2 is devoted to separability, entanglement witnesses, optimality, positive maps, and so on. Then come our main results: In Sec. III.1 we introduce sufficient criteria for optimality of bipartite entanglement witnesses. These criteria are derived from the spanning property; the core result here will be that a witness $W$ has the spanning property if and only if it vanishes on a full-rank separable state. Again using Choi-Jamiołkowski, this translates to a new characterization of the spanning property involving entanglement-breaking channels (Thm. 1), so simply evaluating this on the particular entanglement-breaking channel $X\mapsto\frac{1}{\dim\mathcal{H}+1}(X+{\rm tr}(X){\bf 1})$ leads to a sufficient criterion for optimality where one only has to check the kernel of one of two specific bipartite states (Thm. 2). This in turn implies an even simpler criterion based on a new lower bound on the expectation value of a given witness on any maximally entangled state which, if saturated, guarantees optimality (Coro. 2). Both new criteria of ours are a lot easier to use and test for than the spanning property—analytically as well as numerically—and yet the only witness we could find which (i) has the spanning property but (ii) which our criterion does not detect as optimal is the flip in odd dimensions (Rem. 3).
@@ II.2 Entanglement witnesses & positive maps
Figure 2: Overview of optimality conditions on entanglement witnesses $W\in\mathbb{C}^{m\times m}\otimes\mathbb{C}^{n\times n}$ where, without loss of generality, $m\leq n$. Green boxes are new results proven in this work. (a) Proof of Coro. 2, (b) Proof of Thm. 2, (c) Rem. 3, (d) & (e) Thm. 1, (f) & (h) Lewenstein et al. [16], (g) Augusiak et al. [30, 31], see also Choi et al. [36], (i) Gühne et al. [2], (j) $W=\mathbb{F}+|00\rangle\langle 00|$, because $W-|00\rangle\langle 00|$ is a witness (hence $W$ is not optimal), but $\langle 01|W|01\rangle=0$ (weakly optimal), (k) Coro. 3, (l) $W=\mathbb{F}+|00\rangle\langle 00|$ is weakly optimal but 00 is not an eigenvalue of $W+{\rm tr}_{2}(W)\otimes{\bf 1}$
@@ III.1 New sufficient optimality criteria
Before checking more witnesses we make the following observation: in both previous examples $\ker(W+{\rm tr}_{2}(W)\otimes{\bf 1})$ contained not just a vector of full Schmidt rank, but even a maximally entangled state. It turns out that in this case one finds an even simpler sufficient criterion for optimality via expectation values:
@@
The bound is tight and can be used to easily show optimality of not just the transpose in even dimensions or the reduction witness (cf. also (12) below), but more generally of the Breuer-Hall witness $W_{\rm BH}\in\mathbb{C}^{2n\times 2n}\otimes\mathbb{C}^{2n\times 2n}$ (recall Eq. (7))
@@ III.2 Implications for positive maps and spectra of witnesses
Next, Coro. 2 places a lower bound on $\langle\Omega|\mathsf{C}(\Phi^{\dagger})|\Omega\rangle$ which—if $m=n$—relates to (a version of) the superoperator trace ${\rm tr}(\Phi)=\sum_{j,k=1}^{n}\langle j|\Phi(|j\rangle\langle k|)|k\rangle$ of $\Phi$ [59, Eq. (6)]. Notably, this trace connects to the entanglement fidelity $F_{e}(\Phi,\rho)$ [60], [20, Def. 3.30] in that $F_{e}(\Phi,\rho)={\rm tr}(\Phi(\rho(\cdot)\rho))$ for all $\Phi$ completely positive and all $\rho\geq 0$ (if $\rho$ is the maximally mixed state then this reduces to the channel fidelity $\frac{1}{n^{2}}{\rm tr}(\Phi)$ [61]). The known bound for this quantity from spectral considerations is ${\rm tr}(\Phi)\in[-n^{2}\|\Phi^{\dagger}({\bf 1})\|_{\infty},n^{2}\|\Phi^{\dagger}({\bf 1})\|_{\infty}]$ [62, Coro. 2.3.8], where the upper bound is tight. The lower bound, however, can be strengthened further using our new results:
@@
(i): Because $\Phi$ is positive, $\mathsf{C}(\Phi)$ is block-positive so Coro. 2 implies $\frac{1}{n}\langle\Gamma|\mathsf{C}(\Phi)|\Gamma\rangle\geq-\frac{1}{n}{\rm tr}(\mathsf{C}(\Phi))$. Now $\langle\Gamma|\mathsf{C}(\Phi)|\Gamma\rangle={\rm tr}(\Phi)$ [59, Lemma 2] as well as ${\rm tr}(\mathsf{C}(\Phi))={\rm tr}(\Phi({\bf 1}))={\rm tr}(\Phi^{\dagger}({\bf 1}))$ which combines to the desired inequalities. Finally, equality is achieved, e.g., by the reduction map $R={\rm tr}(\cdot){\bf 1}-(\cdot)$: because $R({\bf 1})=(n-1){\bf 1}$ one has ${\rm tr}(R({\bf 1}))=n^{2}-n$ and hence
@@
(iv): As before the key here is [59, Lemma 2] which states that ${\rm tr}(\Phi(U^{\dagger}(\cdot)U))=\langle{\rm vec}(U)|\mathsf{C}(\Phi)|{\rm vec}(U)\rangle$ for all $U$. Now if $U$ is unitary, then $|\Omega\rangle:=\frac{1}{\sqrt{n}}|{\rm vec}(U)\rangle=\frac{1}{\sqrt{n}}({\bf 1}\otimes U)|\Gamma\rangle$ is maximally entangled so we know that
@@
With this in mind let us revisit the positive maps behind the witnesses we looked at in Sec. III.1. The transpose always leads to optimal witnesses, but satisfies the trace condition from Thm. 3 (iv) only for $n$ even (which is in line with Remark 3 where we saw that in odd dimensions, our sufficient criterion fails for the transpose): ${\rm tr}(U^{\dagger}(\cdot)^{T}U)={\rm tr}(\overline{U}U)$, the minimum of which is either $-n$ if $n$ is even, or $-(n-2)$ if $n$ is odd, cf. Lemma 3 in Appendix B.
The re-scaled reduction map (by $\frac{1}{n-1}$) as well as the re-scaled Breuer-Hall maps (by $\frac{1}{2n-2}$, with $U$ unitary and anti-symmetric) are positive and trace-preserving and they both have trace $-n$, meaning Thm. 3 (iv) implies optimality of the associated witnesses. This optimality is also reflected in the region of positive, trace-preserving maps from [63, Eq.(72) & Lemma 24] which also takes the minimal value $-n$ (for $\alpha=\beta=-\frac{1}{n-1}$ in which case $M=\frac{1}{n-1}R$ with $R$ the reduction map from above), regardless of the dimension $n$.
@@
Finally, given that Coro. 2 is essentially a spectral condition on block-positive matrices (since the trace is the sum of all eigenvalues), and Coro. 1 imposes similar restrictions, let us quickly recap some known spectral constraints on block-positive operators $W\in\mathbb{C}^{m\times m}\otimes\mathbb{C}^{n\times n}$ [64, Sec. 2.3] to see how our results compare: $W$ has no more than $(m-1)(n-1)$ negative eigenvalues (even in the decomposable case), one has ${\rm tr}(W)^{2}\geq{\rm tr}(W^{2})$, and if $W$ is decomposable then the smallest eigenvalue of $W$ satisfies $\lambda_{\rm min}(W)\geq-\frac{1}{2}{\rm tr}(W)$. Moreover, there are some lower bounds on the ratio $\frac{\lambda_{\rm min}(W)}{\lambda_{\rm max}(W)}$ between the smallest and the largest eigenvalue of $W$ [65, Coro. 5.5] (which we will not spell out explicitly here). To add to these constraints, our results imply the following new bounds:
@@ Acknowledgements.
We would like to thank Jens Eisert for insightful discussions and for drawing our attention to key references on the separability problem. We are also grateful to Gregory White for valuable feedback on an earlier version of this manuscript. FvE is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 384846402, and supported by the Einstein Foundation (Einstein Research Unit on Quantum Devices) and the MATH+ Cluster of Excellence. SC is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689).

### v2

@@ I Introduction
In this work we tackle both these problems. As such, this paper is organized as follows: In Sec. II.1 we recap some basic notions of channels and general linear maps, and Sec. II.2 is devoted to separability, entanglement witnesses, optimality, positive maps, and so on. Then come our main results: In Sec. III.1 we introduce sufficient criteria for optimality of bipartite entanglement witnesses. These criteria are derived from the spanning property; the core result here will be that a witness $W$ has the spanning property if and only if it vanishes on a full-rank separable state. Again using Choi-Jamiołkowski, this translates to a new characterization of the spanning property involving entanglement-breaking channels (Thm. 1), so simply evaluating this on the particular entanglement-breaking channel $X\mapsto\frac{1}{\dim\mathcal{H}+1}(X+{\rm tr}(X){\bf 1})$ leads to a sufficient criterion for optimality where one only has to check the kernel of one of two specific bipartite states (Thm. 2). This in turn implies an even simpler criterion based on a new lower bound on the expectation value of a given witness on any maximally entangled state which, if saturated, guarantees optimality (Coro. 2). Both new criteria of ours are a lot easier to use and test for than the spanning property—analytically as well as numerically—and yet the only witness we could find which (i) has the spanning property but (ii) which our criterion does not detect as optimal is the flip in odd dimensions (Rem. 3 (i)).
@@ II.2 Entanglement witnesses & positive maps
Figure 2: Overview of optimality conditions on entanglement witnesses $W\in\mathbb{C}^{m\times m}\otimes\mathbb{C}^{n\times n}$ where, without loss of generality, $m\leq n$. Green boxes are new results proven in this work. (a) Proof of Coro. 2, (b) Proof of Thm. 2, (c) Rem. 3 (i), (d) & (e) Thm. 1, (f) & (h) Lewenstein et al. [16], (g) Augusiak et al. [30, 31], see also Choi et al. [36], (i) Gühne et al. [2], (j) $W=\mathbb{F}+|00\rangle\langle 00|$, because $W-|00\rangle\langle 00|$ is a witness (hence $W$ is not optimal), but $\langle 01|W|01\rangle=0$ (weakly optimal), (k) Coro. 3, (l) $W=\mathbb{F}+|00\rangle\langle 00|$ is weakly optimal but 00 is not an eigenvalue of $W+{\rm tr}_{2}(W)\otimes{\bf 1}$
@@ III.1 New sufficient optimality criteria
  • (ii)
    There is an interesting connection between Thm. 2 and the structural physical approximation (SPA) conjecture. Although the SPA conjecture was disproven [58], it originally proposed—given a witness $W$—to study separability of $({\rm id}\otimes\Psi_{D})(W)$ with $\Psi_{D}(X):=(1-p)X+p\,{\rm tr}(X)\frac{\bf 1}{n}$ and $p=p(W)$ is chosen to be the smallest value for which $({\rm id}\otimes\Psi_{D})(W)\geq 0$ [59, 49, 60]; one always has $p(W)\leq\frac{n}{n+1}$ because $\Psi_{D}$ is entanglement-breaking for $p=\frac{n}{n+1}$, recall Sec. II.2. Notably, SPA re-optimizes $p$ witness-by-witness, whereas we simply fix the universal upper bound $p=\frac{n}{n+1}$ when applying $\Psi_{D}$ because we, crucially, only require $({\rm id}\otimes\Psi_{D})(W)$ to be positive semi-definite (not separable). In this framework, Thm. 2 states that if the minimal $p(W)$ for some witness $W$ is indeed $\frac{n}{n+1}$, and if the corresponding kernel vector has maximal Schmidt rank, then $W$ is optimal.
@@
Before checking more witnesses we make the following observation: in both previous examples $\ker(W+{\rm tr}_{2}(W)\otimes{\bf 1})$ contained not just a vector of full Schmidt rank, but even a maximally entangled state. It turns out that in this case one finds an even simpler sufficient criterion for optimality via expectation values 111 A similar condition could be formulated for the general case of a kernel vector of full Schmidt rank ${\rm vec}(X)$; however, one would have to replace ${\rm tr}(W)$ in (11) by something like ${\rm tr}(({\bf 1}\otimes XX^{\dagger})W)$, thus leading to a state-dependent bound which is not as nice to work with anymore. :
@@
The bound from Coro. 2 is tight and, in particular, equality in (11) for some state $\Omega$ necessitates that $-\,\frac{{\rm tr}(W)}{\min\{m,n\}}$ is an eigenvalue of $W$. Once that is certified, it suffices to check whether the corresponding eigenvector (resp. whether some eigenvector from the corresponding eigenspace) is maximally entangled.
Now Coro. 2 can be used to easily show optimality of not just the transpose in even dimensions or the reduction witness (cf. also (12) below), but more generally of the Breuer-Hall witness $W_{\rm BH}\in\mathbb{C}^{2n\times 2n}\otimes\mathbb{C}^{2n\times 2n}$ (recall Eq. (7))
@@ III.2 Implications for positive maps and spectra of witnesses
Next, Coro. 2 places a lower bound on $\langle\Omega|\mathsf{C}(\Phi^{\dagger})|\Omega\rangle$ which—if $m=n$—relates to (a version of) the superoperator trace ${\rm tr}(\Phi)=\sum_{j,k=1}^{n}\langle j|\Phi(|j\rangle\langle k|)|k\rangle$ of $\Phi$ [63, Eq. (6)]. Notably, this trace connects to the entanglement fidelity $F_{e}(\Phi,\rho)$ [64], [20, Def. 3.30] in that $F_{e}(\Phi,\rho)={\rm tr}(\Phi(\rho(\cdot)\rho))$ for all $\Phi$ completely positive and all $\rho\geq 0$ (if $\rho$ is the maximally mixed state then this reduces to the channel fidelity $\frac{1}{n^{2}}{\rm tr}(\Phi)$ [65]). The known bound for this quantity from spectral considerations is ${\rm tr}(\Phi)\in[-n^{2}\|\Phi^{\dagger}({\bf 1})\|_{\infty},n^{2}\|\Phi^{\dagger}({\bf 1})\|_{\infty}]$ [66, Coro. 2.3.8], where the upper bound is tight. The lower bound, however, can be strengthened further using our new results:
@@
(i): Because $\Phi$ is positive, $\mathsf{C}(\Phi)$ is block-positive so Coro. 2 implies $\frac{1}{n}\langle\Gamma|\mathsf{C}(\Phi)|\Gamma\rangle\geq-\frac{1}{n}{\rm tr}(\mathsf{C}(\Phi))$. Now $\langle\Gamma|\mathsf{C}(\Phi)|\Gamma\rangle={\rm tr}(\Phi)$ [63, Lemma 2] as well as ${\rm tr}(\mathsf{C}(\Phi))={\rm tr}(\Phi({\bf 1}))={\rm tr}(\Phi^{\dagger}({\bf 1}))$ which combines to the desired inequalities. Finally, equality is achieved, e.g., by the reduction map $R={\rm tr}(\cdot){\bf 1}-(\cdot)$: because $R({\bf 1})=(n-1){\bf 1}$ one has ${\rm tr}(R({\bf 1}))=n^{2}-n$ and hence
@@
(iv): As before the key here is [63, Lemma 2] which states that ${\rm tr}(\Phi(U^{\dagger}(\cdot)U))=\langle{\rm vec}(U)|\mathsf{C}(\Phi)|{\rm vec}(U)\rangle$ for all $U$. Now if $U$ is unitary, then $|\Omega\rangle:=\frac{1}{\sqrt{n}}|{\rm vec}(U)\rangle=\frac{1}{\sqrt{n}}({\bf 1}\otimes U)|\Gamma\rangle$ is maximally entangled so we know that
@@
With this in mind let us revisit the positive maps behind the witnesses we looked at in Sec. III.1. The transpose always leads to optimal witnesses, but satisfies the trace condition from Thm. 3 (iv) only for $n$ even (which is in line with Remark 3 (i) where we saw that in odd dimensions, our sufficient criterion fails for the transpose): ${\rm tr}(U^{\dagger}(\cdot)^{T}U)={\rm tr}(\overline{U}U)$, the minimum of which is either $-n$ if $n$ is even, or $-(n-2)$ if $n$ is odd, cf. Lemma 3 in Appendix B.
The re-scaled reduction map (by $\frac{1}{n-1}$) as well as the re-scaled Breuer-Hall maps (by $\frac{1}{2n-2}$, with $U$ unitary and anti-symmetric) are positive and trace-preserving and they both have trace $-n$, meaning Thm. 3 (iv) implies optimality of the associated witnesses. This optimality is also reflected in the region of positive, trace-preserving maps from [67, Eq.(72) & Lemma 24] which also takes the minimal value $-n$ (for $\alpha=\beta=-\frac{1}{n-1}$ in which case $M=\frac{1}{n-1}R$ with $R$ the reduction map from above), regardless of the dimension $n$.
@@
Finally, given that Coro. 2 is essentially a spectral condition on block-positive matrices (since the trace is the sum of all eigenvalues), and Coro. 1 imposes similar restrictions, let us quickly recap some known spectral constraints on block-positive operators $W\in\mathbb{C}^{m\times m}\otimes\mathbb{C}^{n\times n}$ [68, Sec. 2.3] to see how our results compare: $W$ has no more than $(m-1)(n-1)$ negative eigenvalues (even in the decomposable case), one has ${\rm tr}(W)^{2}\geq{\rm tr}(W^{2})$, and if $W$ is decomposable then the smallest eigenvalue of $W$ satisfies $\lambda_{\rm min}(W)\geq-\frac{1}{2}{\rm tr}(W)$. Moreover, there are some lower bounds on the ratio $\frac{\lambda_{\rm min}(W)}{\lambda_{\rm max}(W)}$ between the smallest and the largest eigenvalue of $W$ [69, Coro. 5.5] (which we will not spell out explicitly here). To add to these constraints, our results imply the following new bounds:
@@ Acknowledgements.
We would like to thank Jens Eisert for insightful discussions and for drawing our attention to key references on the separability problem. We also thank Joonwoo Bae making us aware of the connection between our results and the SPA conjecture (cf. Rem. 3 (ii)). We are also grateful to Gregory White for valuable feedback on an earlier version of this manuscript. FvE is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 384846402, and supported by the Einstein Foundation (Einstein Research Unit on Quantum Devices) and the MATH+ Cluster of Excellence. SC is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689).